38 t39 t4D +41 t107
38 t39 t4D +41 t107
This Question: 1 pt Use the following pattern to complete parts (a) and (b) below 18 54 18.2 18 54 9018.32 1854 90+ 126 18.4 a. Conjecture a generalization based on this pattern. Choose the correct answer below OA. The sum of the terms is 18 2 O B. The sum of the terms is 18(n+ 1) O C. The sum of the terms is 18n. 2 O D. The sum of the terms is 18n b. Based on the generalization in (a), find the sum of the sequence 18+54+90+ 126 +234 The sum of the sequence is Click to select your answer(s). Save for Later Type here to search
Answer #1

ahivewi- 18(+2)-182 18 (1t 3) = 18.2- 2 18+S48.22 2. 18 (+ 3+S) 18 3 2 1 ((2-1)(21)-1) -+ (23)-))) .&81 2 1tt y 90+ 126 = 18.